OpenStudy (gabylovesyou):

@Night-Watcher The theme park company is building a scale model of the killer whale stadium main show tank for an investor's presentation. Each dimension will be made 6 times smaller to accommodate the mock-up in the presentation room. How many times smaller than the actual volume is the volume of the mock-up?

OpenStudy (blackops2luvr):

SF = AF^2 = VF^3

OpenStudy (blackops2luvr):

\(\Huge\cal\color{lime}{:P}\)

OpenStudy (blackops2luvr):

did that help you at all?

OpenStudy (blackops2luvr):

That means scale factor (1d) = area factor (2d) ^2 = volume factor (3d) ^3

OpenStudy (anonymous):

Well yes that the formula i was about to provide. But @blackops2luvr got it in first! >.< lol

OpenStudy (blackops2luvr):

So if each dimension, which is a 1d measurement is made 6 times smaller, so 1/6 of the size, the volume will have changed by that ^3'd. So 1/6x1/6x1/6 which = 1/216 of the volume

OpenStudy (blackops2luvr):

lol, im only in 8th XD

OpenStudy (blackops2luvr):

it's 216 times smaller, so that is 1/216 of the original volume. Multiply that by 100 to get 0.4629% of the volume

OpenStudy (blackops2luvr):

do you understand?

OpenStudy (gabylovesyou):

kinda... im going to read again .-.

OpenStudy (blackops2luvr):

lol ok

OpenStudy (gabylovesyou):

thats u ?

OpenStudy (blackops2luvr):

oops wrong one >.< im an idiot

OpenStudy (blackops2luvr):

well yea

OpenStudy (gabylovesyou):

rofl cx mk

OpenStudy (blackops2luvr):

\(\small\cal\color{red}{R}\)\(\small\cal\color{orange}{a}\)\(\small\cal\color{yellow}{i}\)\(\small\cal\color{green}{n}\)\(\small\cal\color{blue}{b}\)\(\small\cal\color{purple}{o}\)\(\small\cal\color{red}{w}\) POWA

OpenStudy (gabylovesyou):

cx

OpenStudy (blackops2luvr):

:3 i live in florida

OpenStudy (gabylovesyou):

o.o me too

OpenStudy (blackops2luvr):

XD but im 13 >.< do u have kik?

OpenStudy (gabylovesyou):

i pmed it.. im 15 cx

OpenStudy (gabylovesyou):

im going to orlando this weekend ;o

OpenStudy (blackops2luvr):

XD

OpenStudy (blackops2luvr):

do u watch South beach tow? c;

OpenStudy (gabylovesyou):

nope.. never heard of it cx ima kik u now..

OpenStudy (blackops2luvr):

i cant now XD my tab is dead, maby around 12:00 XD

OpenStudy (blackops2luvr):

thts when i watch my movies but ill kik ya XD

OpenStudy (gabylovesyou):

ight cx

OpenStudy (blackops2luvr):

oh 12 am not pm btw

OpenStudy (blackops2luvr):

Rigby <3

OpenStudy (blackops2luvr):

Well, i gtg D; bank.. cya

OpenStudy (anonymous):

I came up with 28984.44581%

OpenStudy (gabylovesyou):

how

OpenStudy (gabylovesyou):

nevermind i know how